OpenStudy (anonymous):

A company estimates that f(x) thousand software games can be sold at the price of $x as given in the table. x 20 80 100 f(x) 84 120 168 Estimate the number of games that can be sold at: a) $24 b) $36

OpenStudy (anonymous):

@Twis7ed can u help?

OpenStudy (anonymous):

You could use linear interpolation in order to find out the slope of the curve if it was a straight line from say x = 20 to x = 80 and then, using the slope of that line estimate that value of f(x).

OpenStudy (anonymous):

can you show it how i will do it?

OpenStudy (anonymous):

Ok, so you want to get a reasonably good approximation of the slope of the curve at points x = 24 and x = 36. To do this you get the data that is closest to those intervals that would be, in this case, from x = 20 to x = 80. Using that you can figure out that the slope between these two points would be \[\frac{ 120-84 }{ 80-20 }\] which means a slope of 0.6. Using that, just solve for what the value of the function would be at points 24 and 36. You won't get an exact answer since 0.6 is not the exact slope of the curve in those intervals but you get a decent approximation.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so how would i solve for those values?

OpenStudy (anonymous):

Well, you would know the slope between each x-value so you would just multiply the slope (0.6 in this case) by the change in x from 20 to the value you need and add 84. ^Not sure if that explanation makes much sense :P

OpenStudy (anonymous):

lol okay thanks

OpenStudy (anonymous):

Hey @Twis7ed i tried to multiply 0.6 into x 20 into 84 but i think i idid it wrong so can you show it ? thanks

OpenStudy (anonymous):

What did you get as your answer?

OpenStudy (anonymous):

80 :/

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

No, I can't seem to explain it well so I'll just give you a link with an explanation of the process, http://www.asz.ymmf.hu/talata/math/Exercises/ch03_01.pdf go to page 246 on that and it explains how to do it in a similar setting.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so i got L(x) = f(0.6) + f'(0.6) (x - 0.6) is that correct?

OpenStudy (anonymous):

like plugging it in

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

okay so now i just solve it right?

OpenStudy (anonymous):

Yeah, just plug in the values for 24 and 36 and you'll get the answer

OpenStudy (anonymous):

plug those values where?

OpenStudy (anonymous):

For a) \[L(24) = 80+0.6(24-20)\] and then for b) all you would have to do is change the value for x

OpenStudy (anonymous):

I got f(0.6) + f(0) ( x - 0.6) after i solved

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

so L(36) = 20 + 0.6 (36 - 20) ?

OpenStudy (anonymous):

Almost, remember the f(20) is 80 so it would be L(36)=80+0.6(36-20)

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

okay so then i just solve

OpenStudy (anonymous):

Yep :D

OpenStudy (anonymous):

thanks :)